dentifying the Line of Best Fit (Least Squares Regression)

Question

A random sample of 11 workers produced the following data, where x is the years of employment, and y is the average number of minutes required to complete the morning duties. The data are presented below in the table of values.

x y
5 49
8 42
10 41
12 39
15 37
16 34
19 30
22 25
25 24
27 22
28 16

What is the equation of the regression line?

Select the correct answer below:

yˆ=10.082x+54.2

yˆ=−1.267x+17.0

yˆ=−1.267x+54.2

yˆ=10.082x+17.0
Answer #1
X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
5 49 144.00 267.8 -196.4
8 42 81.00 87.7 -84.3
10 41 49.00 70.0 -58.5
12 39 25.00 40.5 -31.8
15 37 4.00 19.0 -8.7
16 34 1.00 1.9 -1.4
19 30 4.0 7.0 -5.3
22 25 25.00 58.3 -38.2
25 24 64.00 74.6 -69.1
27 22 100.00 113.1 -106.4
28 16 121.00 276.77 -183.00
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 187 359 618.000 1016.5 -783
mean 17.00 32.64 SSxx SSyy SSxy

sample size ,   n =   11          
here, x̅ =   17.00   ,   ȳ =   32.6363
                  
SSxx =    Σ(x-x̅)² =    618.00          
SSxy=   Σ(x-x̅)(y-ȳ) =   -783.0          
                  
slope ,    ß1 = SSxy/SSxx =   -1.2670          
                  
intercept,   ß0 = y̅-ß1* x̄ =   54.1752          
                  
so, regression line is   Ŷ =   −1.267x + 54.2